OpenStudy (anonymous):

a right triangle contains a 28degree angle whose adjacent side measures 10 cm. what is the length of the hypontenus?

OpenStudy (anonymous):

Use cosine ratio to find it. cos 28 = 10 / hypontenus hypontenus = 10/ (cos28) = ...?

OpenStudy (anonymous):

howd you do that?

OpenStudy (anonymous):



OpenStudy (anonymous):

/ means multiplyy or divide?

OpenStudy (anonymous):

You need a trig calculator or a table. Set it up as\[\cos(28^{o})=\frac{10}{h}\implies h=\frac{10}{\cos(28^{o})}\]

OpenStudy (anonymous):

/ = divide